Write an equation for the line that is parallel to the given line and contains the given point. Write your answer in standard form with integer coefficients. 2x+3y=30; (2,-5)
Answer by mananth(16946)

2 x + 3 y = 30
Find the slope of this line
make y the subject
3 y = -2 x + 30
Divide by 3
y = -2/3 x + 10
Compare this equation with y=mx+b
slope m = -2/3
The slope of a line parallel to the above line will be the same
The slope of the required line will be -2/3
m= -2/3 ,point ( 2 , -5 )
Find b by plugging the values of m & the point in
y=mx+b
-5 = -4/3 + b
b= -11/3

m= -2/3
Plug value of the slope and b in y = mx +b
y = - 2/ 3 x -11/ 3
multiply by 3
3y=-2x-11
2x+3y=-11
m.ananth@hotmail.ca